A closed vessel contains 2 kg of carbon-dioxide at temperature 20 oC and pressure 0.7 bar. Heat is supplied to the vessel till the gas occupies a pressure of 14 bar, Calculate (a) final temperature (h) work done on or by the gas, (e) heat added and (d) change in internal energy e) Change in entropy. Take specific heat of gas at constant volume as 0.657 kJ/kg K. 

Grade:12

1 Answers

Arun
6 years ago
Dear Manish
 
for constant volume process
P1/T1 = P2/T2
T2 = 1.4*293 / 0.7
T2 = 586 K
hence
Final temperature = 586 K
work done W = p * dV = 0
Heat added = m * Cp (T2 – T1) = 2 * 0.657 * 293 = 385 KJ
Change in internal energy U12 = Q12 – W12 = 385 – 0 = 385 KJ
 
Regards
Arun (askIITians forum expert)